OpenStudy (j-cross18):

Who created the theory 2+2=5?

OpenStudy (zale101):

this theory doesn't exist

OpenStudy (rina.r):

theory of what?

OpenStudy (j-cross18):

Like how does it work?

OpenStudy (ashleyisakitty):

2+2=4

OpenStudy (j-cross18):

@Rina.r what do you mean by that?

OpenStudy (anonymous):

2 + 2 does not = 5

OpenStudy (anonymous):

\[\frac{ 2 }{ 0 }+\frac{ 2 }{ 0 }=\frac{ 5 }{ 0 }\]

OpenStudy (j-cross18):

I know that but why does that keep popping up as the correct answer on my test for algebra 2 like it would say 2+2=5 true or false and i say false and get it wrong. Then i go to ask my teacher how did i get it wrong and she said because someone made a theory that 2+2=5 and she wants me to go find who so now im asking who?

OpenStudy (anonymous):

Did you have a lesson to do on it, if so look in the lesson it might tell you.

OpenStudy (j-cross18):

No i didnt unfortunetly i did not so does anyone know who came up with that theory???

OpenStudy (anonymous):

that is kind of stupid to test you on something without teaching it do you work online if so what class and through what online program

OpenStudy (j-cross18):

im not getting your question @Victoria_Davis_2018

OpenStudy (anonymous):

Literally the only two things that I found was that it was used as a joke in one sense: "If you do 2.3 rounded (2) + 2.4 rounded (2), you get 4.7 which rounds to 5." And the other had to do with a quote by George Orwell from a book. ...But with that being said I have no idea what that has to do with your test lol.